OpenStudy (anonymous):

A race car's velocity decreases from 36 m/s to 15 m/s over a 3.0 s period, what is the average acceleration? A race car's velocity decreases from 36 m/s to 15 m/s over a 3.0 s period, what is the average acceleration? @Physics

OpenStudy (mattfeury):

average acceleration = change in velocity / change in time.

OpenStudy (anonymous):

I did the problem and got -63m/s(squared). Is this correct?

OpenStudy (mattfeury):

\[\frac{v_f - v_1}{\Delta t} = \frac{15 - 36}{3} = -7\]

OpenStudy (stormfire1):

\[\frac{Vfinal-Vinitial}{t}=-7m/s^2\]\[\frac{15m/s-36m/s}{3.0s}=-7m/s^2\]